OpenStudy (anonymous):

solve for p s=p/3 + 7q

OpenStudy (anonymous):

Subtract 7q from both sides. p/3 = s - 7q. then multiply by 3. p = 3s - 21q.

OpenStudy (anonymous):

that is the final answer? p= 3s-21q